# RobotFingerPrint: Unified Gripper Coordinate Space for Multi-Gripper Grasp Synthesis

### v1

@@ III-C Model Architecture
Figure 6: Method Overview: A randomly sampled latent vector is decoded to a probable object coordinate map $\hat{\Phi}_{O}$. Given a target (possibly unseen) gripper, correspondence is established between object and gripper points using $\hat{\Phi}_{O}$. The same $\hat{\Phi}_{O}$ also guides the pose initialization for the grasp optimization which results in successful grasps.
@@
The overall goal for our RFP method is to synthesize grasps on objects and generalize to unseen grippers and objects. In our modeling setup, we assume that we are given an object point cloud $P_{O}$ with $N$ points. The prediction problem consists of predicting the intermediate grasp representation of object coordinate map $\Phi_{O}$ with a values $(\lambda,\varphi)$ for each of the $N$ points. We pose this as a generative modeling problem and utilize the CVAE architecture used in previous works on hand-object grasping [25, 5].
@@
The goal for the learning model is to reason about the grasps on the object $O$ and predict the intermediate grasp representations in the form of object coordinate maps $\Phi_{O}$. Here each grasp $q$ on $O$ is represented by its associated coordinate map $\Phi_{O}$. Given $P_{O}$ and its coordinate map $\Phi_{O}$, the encoder $p_{\alpha}$ of the C-VAE model learns to encode the object point cloud and the coordinate map into a latent vector $z$ by modeling the distribution $p_{\alpha}(z|P_{O},\Phi_{O})$. The decoder $p_{\beta}$ learns to predict an accurate $\hat{\Phi}_{O}$ on the object given a randomly sampled $z$ by modeling the distribution $p_{\beta}(\Phi_{O}|z,P_{O})$. We utilize PointNet++ [26] based backbone, where the final output consists of $(\hat{\lambda},\hat{\varphi})$ coordinate prediction for each point in the object point cloud.
@@
The model is learned by using a standard loss formulation for CVAE with two components for the training objective function. The first component consists of the reconstruction loss $L_{recon}$ on the ground truth coordinate map $\Phi$ and its prediction $\hat{\Phi}$: $L_{recon}=\sum_{i=1}^{N}||\Phi_{O}^{(i)}-\hat{\Phi}_{O}^{(i)}||$. The second component consists of a KL-divergence term $L_{KL}$ between a prior distribution on the latent space $p_{z}=\mathcal{N}(0,1)$ and the approximate posterior modeled by the decoder $p_{\beta}(\Phi_{O}|z,P_{O})$: $L_{KL}=D_{KL}(p_{z},p_{\beta})$. The total loss is a weighted sum between the two: $L_{total}=w_{1}\;L_{recon}\;+\;w_{2}\;L_{KL}$. During inference, given an object point cloud, we sample a latent vector $z\sim\mathcal{N}(0,1)$ and pass it through the decoder along with the point cloud to obtain the predicted coordinate map $\hat{\Phi}_{O}$. This output is then used as an intermediate representation for grasp optimization on a given gripper as shown in Fig. 6.
III-D Grasp Optimization
III-D1 Coordinate Map Correspondence
@@
The ultimate goal of the method is to produce a grasp configuration $q_{G}$ with pose and joint values on an object $O$ and a target gripper $G$. We generate $q_{G}$ for a predicted object coordinate map $\hat{\Phi}_{O}$ and $G$ via a correspondence-driven optimization with the gripper coordinate map $\Phi_{G}$ on its in interior surface point set $P_{G}$. In contact map based grasp synthesis, the predicted contact map is oblivious to the gripper fingers and hence ambiguous to interpret. A key advantage of UGCS is that we can easily establish a correspondence between an object $O$ with $\hat{\Phi}_{O}$ and gripper $G$ with $\Phi_{G}$. Given $N$ object points, and $M$ gripper points, we establish a correspondence between them by assigning each object point $v_{o}$ to a gripper point $v_{g}^{*(o)}$ based on the great-circle distance $D_{H}$ (also known as the Haversine distance) between their respective coordinates $(\hat{\lambda}_{v_{o}},\hat{\varphi}_{v_{o}})$ and $(\lambda_{v_{g}},\varphi_{v_{g}})$. The distance is computed after projecting the coordinates back from the normalized space to the original ranges for latitude and longitude.
$v_{g}^{*(o)}=\operatorname*{\arg\!\min}_{v_{g}\in P_{G}}\;\;D_{H}\left((\hat{\lambda}_{v_{o}},\hat{\varphi}_{v_{o}})\;;\;(\lambda_{v_{g}},\varphi_{v_{g}})\right).$
Using this procedure we obtain a dense mapping between the object and gripper points and it can guide the grasp optimization towards suitable grasps for $\hat{\Phi}_{O}$. While establishing this correspondence, we do not consider probable no-contact object points which are inferred on the basis of low values for both of their predicted coordinates. The final result from the correspondence are two matching subsets $P_{O_{c}}\subset O$ and $P_{G_{c}}\subset P_{G}$ from the object and gripper points, respectively.
III-D2 Pose Initialization
@@
A possible way to initialize the optimization procedure for $q_{G}$ would be to sample random gripper poses around the object via some heuristics. However, this can lead to suboptimal grasps due to improper initialization. We observe that in addition to the correspondence, the predicted coordinate map $\Phi_{O}$ contains another signal to the optimization procedure in terms of probable gripper pose. During the construction for UGCS, each of the grippers grasps the sphere with their palm touching the sphere’s south pole with a normalized coordinate value $(\lambda=0,\varphi=1)$. As a consequence, on the predicted coordinate map $\hat{\Phi}$, we infer a set $S_{B}$ of probable object points that map to gripper base/palm by thresholding on $\hat{\lambda}$ and $\hat{\varphi}$ using pre-defined thresholds $\lambda_{ub}$ and $\varphi_{lb}$, respectively:
$S_{B}=\{v_{o}\in P_{O}\;s.t\;\hat{\lambda}_{v_{o}}\leq\lambda_{ub}\;\;\text{and}\;\;\hat{\varphi}_{v_{o}}\geq\varphi_{lb}\}.$
@@
The initial gripper position is set to $0.1m$ away from the mean of palm-matching set $S_{B}$ along the mean surface normal for it. The initial gripper 6D pose is determined by aligning the gripper palm normal to the opposite of the mean surface normal from $S_{B}$ as shown in Fig. 6.
@@ III-D3 Optimization
The grasp configuration $q_{G}$ is obtained by minimizing an objective function that considers hand-object penetration, joint value validity and hand-object distance. A differentiable kinematics model is used to provide gradients on $q$ using the objective function. Note that the hand-object distance is only computed with the corresponding point pairs from $P_{O_{c}}$ and $P_{G_{c}}$. The optimization problem is defined as
$q_{G}^{*}=\operatorname*{\arg\!\min}_{q}\;E_{dist}(P_{O_{c}},P_{G_{c}};q)+E_{pen}(O,G;q)+E_{n}(q).$
@@
The terms for penetration $E_{pen}$ and joint value validity $E_{n}$ are adopted from [5]. They constrain the grasp $q$ to not collide with the object, and keep the joint values within specified limits. The correspondence-driven $E_{dist}$ acts as a denser version of Inverse Kinematics (IK) by not being limited to only a pre-specified set of sparse keypoints.
IV Experiments
@@
The grasps generated by our method were evaluated using the same simulation protocol used by [5, 12]. Each grasp is tested for 1 second with 60 simulation steps by applying a consistent acceleration of $0.5ms^{-2}$ on the object in the NVIDIA Isaac Gym simulation. This process is repeated six times across the positive and negatives axes of the world coordinate frame. A grasp is considered as a failure if the object moves by 2cm at the end of any test. A grasp is labeled as successful if it passes all six tests. We report the same evaluation metrics of grasp success rate and the diversity in successful grasps (standard deviation of the joint values) in order to be consistent with the comparison. The test set comprised of the same split of 10 unseen objects as in [5, 12], and we sampled 64 grasps for each gripper-object pair for evaluation.
IV-A Grasp Synthesis with Seen Grippers
@@
First, we evaluate the RFP by training with all 5 grippers and evaluating on the test object set. This experiment setting is same as in [12]. We report metrics for Ezgripper, Barrett and Shadowhand as these three grippers have their test results from [5]. Table I shows the success rate and diversity of all the methods using models trained on all the grippers. Results for DFC [19], AdaGrasp [23], GenDexGrasp [5] and GeoMatch [12] are taken from GeoMatch [12]. We observe that RFP outperforms comparable methods in terms of grasp diversity and success rate. Our results are comparable to [12] and better than [5]. It is important to note that while the RFP demonstrates competitive performance with [12], it is capable of dealing with unseen grippers and can generalize well without explicit retraining. Fig. 7 shows some successful grasps from our method.
@@
GDG 77.19 0.207
Figure 7: Successful grasps on unseen objects with RFP.
IV-B Grasp Synthesis with Unseen Grippers
@@
IV-C Ablation Study: Grasp Refinement in Simulation
@@
We carried out an ablation study for the grasp refinement step used in the simulation evaluation tests as in [5, 12]. A contact-aware refinement is applied to the predicted grasp $q$ via a single gradient update step to obtain force closure around the object. By removing this refinement, we wanted to test the efficacy of our method in terms of its original output. We compare it with the refined outputs from other methods under both in-domain and out-domain settings for the grippers. As seen in Tables I and Table II, RFP without refinement remains competitive even under the removal of this refinement step and establishes strong baselines. It even outperforms on some metrics from prior works [5, 19, 23] which include this refinement during their evaluation. The dense correspondence-driven grasp optimization and a stable pose initialization are able to help the optimization, avoiding bad minima and converging to good quality grasps.
Figure 9: Failures from the simulation test due to insufficient force closure or not utilizing all the fingers for grasping.
IV-D Ablation Study: Same Prediction for Different Grippers
@@
Another possible reason for the competitive performance with unseen grippers is visualized in Fig. 8. On the same predicted coordinate map on the object, we observe that the three different grippers grasp the object in similar fashion, especially in terms of approach and position around the object. The mean difference in grasp positions was $0.03m$ for Barrett and Ezgripper while it was around $0.06m$ for both of them with the Shadowhand grasps. This indicates the efficacy of RFP in generalizing the grasps from different grippers in a unified coordinate space. We also present some examples of failure cases in Fig. 9.
V Conclusion and Future Work

### v2

@@ IV-A Model Architecture
Figure 6: Method Overview: A randomly sampled latent vector is decoded to a probable object coordinate map $\hat{\Phi}_{O}$. Given a target (possibly unseen) gripper, correspondence is established between object and gripper points using $\hat{\Phi}_{O}$. $\hat{\Phi}_{O}$ also guides the pose initialization for the grasp optimization which results in successful grasps.
@@
One of the goals for RFP is to synthesize grasps on in/out of domain objects and grippers. In our modeling setup, we assume that we are given an object point cloud $P_{O}$ with $N$ points. The prediction problem consists of inferring an intermediate grasp representation of the object coordinate map $\Phi_{O}$ with values $(\lambda,\varphi)$ for $N$ points. We pose this as a generative modeling problem and utilize the CVAE architecture used in previous work on hand-object grasping [29, 5].
@@
The goal of the learning model is to reason about the grasps on the object $O$ and predict the grasp representations in the form of object coordinate maps $\Phi_{O}$ for each grasp on it. Given $P_{O}$ and its coordinate map $\Phi_{O}$, the encoder $p_{\alpha}$ of the CVAE model learns to encode the object point cloud and the coordinate map into a latent vector $\mathbf{z}$ by modeling the distribution $p_{\alpha}(\mathbf{z}|P_{O},\Phi_{O})$. The decoder $p_{\beta}$ learns to predict an accurate $\hat{\Phi}_{O}$ on the object given a randomly sampled $\mathbf{z}$ by modeling the distribution $p_{\beta}(\Phi_{O}|\mathbf{z},P_{O})$. We utilize a PointNet++ [30] based backbone, where the final output consists of $(\hat{\lambda},\hat{\varphi})$ coordinate prediction for each point in the object point cloud.
@@
The model is trained using a standard loss formulation for CVAE with two components for the objective function. The first component is the reconstruction loss $L_{\text{recon}}$ on the ground truth coordinate map $\Phi$ and its prediction $\hat{\Phi}$: $L_{\text{recon}}=\sum_{i=1}^{N}||\Phi_{O}^{(i)}-\hat{\Phi}_{O}^{(i)}||$. The second component consists of a KL-divergence term $L_{\text{KL}}$ between a prior distribution on the latent space $p_{z}=\mathcal{N}(0,\mathcal{I})$ and the approximate posterior modeled by the decoder $p_{\beta}(\Phi_{O}|\mathbf{z},P_{O})$: $L_{\text{KL}}=D_{\text{KL}}(p_{z},p_{\beta})$. The total loss is a weighted sum between the two: $L_{\text{total}}=\;L_{\text{recon}}\;+\;w_{\text{KL}}\;L_{\text{KL}}$, where the weight $w_{\text{KL}}$ is annealed over training epochs. During inference, given an object point cloud, we sample a latent vector $\mathbf{z}\sim\mathcal{N}(0,\mathcal{I})$ and pass it through the decoder with the object point cloud to obtain the predicted coordinate map $\hat{\Phi}_{O}$. This output is then used as an intermediate representation for grasp optimization on a given gripper as shown in Fig. 6.
IV-B Grasp Optimization
IV-B1 Coordinate Map Correspondence
@@
The final output for grasp synthesis should be a grasp configuration $\mathbf{q}_{G}$, with gripper pose and joint values. We generate $\mathbf{q}_{G}$ for a predicted object $O$’s coordinate map $\hat{\Phi}_{O}$ and target gripper $G$ via a UGCS correspondence optimization. In contact map based grasp synthesis, predicted contact map is oblivious to the gripper fingers and hence ambiguous to interpret. A key advantage of our method is that we can easily establish a correspondence between object and gripper points via the coordinate maps. Given $N$ object points, and $M$ gripper points, correspondence is found by assigning each object point $\mathbf{v}_{o}$ to a gripper point $\mathbf{v}_{g}^{*(o)}$. Assignment is based on the great-circle distance $D_{H}$ (Haversine distance) between their respective coordinates $(\hat{\lambda}_{\mathbf{v}_{o}},\hat{\varphi}_{\mathbf{v}_{o}})$ and $(\lambda_{\mathbf{v}_{g}},\varphi_{\mathbf{v}_{g}})$:
$\mathbf{v}_{g}^{*(o)}=\operatorname*{\arg\!\min}_{\mathbf{v}_{g}\in P_{G}}\;\;D_{H}\left((\hat{\lambda}_{\mathbf{v}_{o}},\hat{\varphi}_{\mathbf{v}_{o}})\;;\;(\lambda_{\mathbf{v}_{g}},\varphi_{\mathbf{v}_{g}})\right).$ (3)
This establishes a dense mapping with two matching subsets of $P_{O_{c}}$ and $P_{G_{c}}$ from the object and gripper points.
IV-B2 Pose Initialization
@@
Randomly sampling gripper poses around the object for initialization like in prior works can lead to suboptimal grasps as the optimization gets stuck in local minima. In addition to the correspondence, the predicted coordinate map $\hat{\Phi}_{O}$ contains another signal to the optimization in terms of gripper pose initialization. By definition, in UGCS, each gripper grasps its sphere with palm/base touching the sphere’s south pole with a normalized coordinate value $(\lambda=0,\varphi=1)$. Using this fact, we infer a set $S_{B}$ of probable object points that map to the gripper base/palm region. $S_{B}$ contains the points whose predicted coordinates $(\hat{\lambda},\hat{\varphi})\in\hat{\Phi}_{O}$ are within some thresholds $\lambda_{\text{ub}}$ and $\varphi_{\text{lb}}$ for base pole’s coordinates:
$S_{B}=\{\mathbf{v}_{o}\in P_{O}\;s.t.,\;\hat{\lambda}_{\mathbf{v}_{o}}\leq\lambda_{\text{ub}}\;\;\text{and}\;\;\hat{\varphi}_{\mathbf{v}_{o}}\geq\varphi_{\text{lb}}\}.$ (4)
@@
The initial gripper position is set to $0.1m$ away from the mean location along the mean surface normal for palm-matching point set $S_{B}$. The approach direction for the gripper palm is opposite to the surface normal as shown in Fig. 6.
@@ IV-B3 Optimization
The grasp configuration $\mathbf{q}_{G}$ is computed by minimizing an objective function that considers hand-object penetration, joint value validity, and hand-object distance. A differentiable kinematics gripper model provides gradients on $\mathbf{q}$. Here, the hand-object distance $E_{\text{dist}}$ is computed only with corresponding point pairs from $P_{O_{c}}$ and $P_{G_{c}}$.
$\mathbf{q}_{G}^{*}=\operatorname*{\arg\!\min}_{\mathbf{q}}E_{\text{dist}}(P_{O_{c}},P_{G_{c}};\mathbf{q})+E_{\text{p}}({O},G;\mathbf{q})+E_{\text{n}}(\mathbf{q}).$ (5)
@@
The terms for hand-object collision $E_{\text{p}}$ and gripper joint validity $E_{\text{n}}$ are adopted from [5]. They constrain the grasp $\mathbf{q}$ to avoid collisions, and keep the joint values within limits. The correspondence-driven $E_{\text{dist}}$ acts as a denser version of inverse kinematics with sparse keypoints seen in prior works.
Figure 7: Grasp Transfer Pipeline: Human demonstration to executing transferred grasp on Fetch
V Grasp Transfer
The unified gripper coordinate space also enables grasp transfer between any two (potentially unseen) grippers without retargeting. Any new, unseen gripper is fit to UGCS via its maximal sphere and subsequently obtaining its surface point coordinate map. Suppose we are given two grippers, $s$ (source) and $t$ (target) with their coordinate maps $\Phi_{s},\Phi_{t}$. In the grasp transfer problem, we are given a grasp $\mathbf{q}_{s}$ for the source and we need to infer a similar grasp $\mathbf{q}_{t}$ for the target. $\mathbf{q}_{t}$ is inferred via the grasp transfer optimization shown in Eq. (6). Here we do not assume access to object point cloud for the source gripper demonstration and consequently, do not involve hand-object point correspondence and collision terms. The two grasps $\mathbf{q}_{s},\mathbf{q}_{t}$ give us the transformed gripper point clouds $P_{s},P_{t}$ by applying the grasps to respective grippers. A set of mutually corresponding pairs of points $P_{s}^{(c)},P_{t}^{(c)}$ via their respective coordinate maps $\Phi_{t},\Phi_{s}$ is computed using a similar formulation seen in Eq. (3). Given the corresponding set, we enforce a distance similarity between the point sets so that the target gripper points align with the source gripper.
$\mathbf{q}_{t}^{*}=\operatorname*{\arg\!\min}_{\mathbf{q}}\;E_{\text{dist}}(P_{s}^{(c)},P_{t}^{(c)};\mathbf{q})+E_{\text{n}}(\mathbf{q}),$ (6)
where $E_{\text{dist}}$ and $E_{\text{n}}$ are defined as in Eq. (5). Note that going from dexterous grippers to simple, 2-finger gripper is not well defined for plain contact-map driven transfer. The dense correspondences from our method resolve this ambiguity and output a functional and similar grasping pose. An illustration of the grasp transfer pipeline between human hand and Fetch robot gripper is shown in Fig. 7.
VI Experiments
VI-A Grasp Synthesis in Simulation
@@
We used the MultiDex dataset [5] to train and evaluate our models in simulation making it comparable to prior works [5, 12] and their evaluation protocol. Each grasp generated by our method is tested for 6 seconds with 60 simulation steps per second and by applying a consistent acceleration of $0.5ms^{-2}$ on the object in the NVIDIA Isaac Gym simulation. A grasp is considered as failure if the object moves by 2cm at the end of any test else it is labeled as successful. We record the same evaluation metrics of grasp success rate and the diversity in successful grasps (standard deviation of joint values) in order to be consistent with the comparison. The test set was comprised of the same split of 10 unseen objects as in [5, 12], and we sampled 64 grasps for each gripper-object pair for evaluation.
VI-A1 Grasp Synthesis with Seen Grippers
@@
We evaluated the RFP by training with all 5 grippers and evaluating on the test object set. We report metrics for Ezgripper, Barrett and Shadowhand according to [5, 12]. Table I shows the success rate and diversity of all the methods using models trained on all the grippers. Results for DFC [20], AdaGrasp [24], GenDexGrasp [5] and GeoMatch [12] are taken from GeoMatch [12]. We observe that RFP outperforms comparable methods in terms of grasp diversity and success rate. Our results are comparable to [12] and better than [5]. It is important to note that while the RFP demonstrates competitive performance with [12], it is capable of dealing with unseen grippers whereas [12] cannot do so. Fig. 8 shows some successful and failure grasps of our method.
@@
GDG [5] 77.19 0.207
Figure 8: Simulation Results: (a) Successful grasps on unseen objects with RFP ; (b) Some failure modes: insufficient force closure or not using all fingers
VI-A2 Grasp Synthesis with Unseen Grippers
@@
Figure 10: Illustration of our real-world grasping pipeline. Our method is used to synthesize grasps given a 3D object model.
Figure 11: Grasp Transfer Experiments: Examples of huamn demo (from robot’s view) with success/failure cases.
VI-B Ablation Studies
VI-B1 Grasp Refinement during Simulation Test
@@
We carried out an ablation study for the grasp refinement step used in the simulation evaluation tests as in [5, 12]. A contact-aware refinement is applied to the predicted grasp $\mathbf{q}$ via a single gradient update step to obtain force closure around the object. By removing this refinement, we wanted to test the efficacy of our method in terms of its original output. We compare it with the refined outputs from other methods under both in-domain and out-domain settings for the grippers. As seen in Tables I and Table II, RFP without refinement remains competitive even under the removal of this refinement step and establishes strong baselines. It even outperforms on some metrics from prior works [5, 20, 24] which include this refinement during their evaluation. The dense correspondence-driven grasp optimization and a stable pose initialization are able to help the optimization, avoiding bad minima and converging to good quality grasps.
VI-B2 Grasp Inference on Same Object Coordinate Map
@@
On the same predicted coordinate map of the object, we observe that the three different grippers grasp the object in a similar fashion, especially in terms of approach and position around the object as visualized in Fig. 9. The mean difference in grasp positions was $0.03m$ for Barrett and Ezgripper while it was circa $0.06m$ for both of them with the Shadowhand grasps. This indicates the efficacy of RFP in generalizing the grasps from different grippers in a unified coordinate space.
VI-C Real World Grasping Evaluation
We used the SceneReplica [16] real-world benchmark to evaluate the grasps generated by our method. SceneReplica benchmark is comprised of 20 standardized cluttered scenes with 5 YCB objects each. It benchmarks the task of grasping each object in scene and moving to a dropoff location with 100 trials in total. Grasping and pick-place success rates were the evaluation metrics used for the trials. An example setup of the experiments is shown in Fig. 10. For our model-based pipeline, we utilized the GDRNPP model [31] for object pose estimation from the robot’s RGB-D camera. Grasp generation was performed for each object using RFP by inferring coordinate maps on the object using the trained model, and optimizing for the Fetch gripper’s grasp configuration. The MultiGripperGrasp [22] toolkit was used to rank the grasps based on a fall time evaluation metric in simulation. Moveit was used for motion planning to evaluate and execute collision-free grasps in descending order of evaluation scores.
It is worth noting that of these 16 YCB objects, 6 objects for which we are generating grasps are unseen to our model. Moreover, the Fetch robot’s 2 finger gripper is also a new gripper outside of the model’s training domain. The grasp was deemed successful if the robot lifted the object, with the pick-and-place task failing only if the object fell before placement. As seen in Table III, the grasps computed by RFP outperform those given by GraspIt! on the benchmark. Despite the training dataset not containing grasps from an equivalent 2-finger gripper, RFP is able to synthesize reasonable real-world grasps for the Fetch gripper. Please see supplementary video for grasping examples.
TABLE III: Grasping and Pick-and-Place (P&P) success rates on the SceneReplica benchmark [16] for 100 trials.
Perception Module Grasp Module P&P (%) ↑ Grasping (%)↑
GDRNPP [31] RFP (ours) 70 73
GDRNPP [31] GraspIt! [1] 66 69
PoseRBPF [32] GraspIt! [1] 58 64
VI-D Human-to-Robot Grasp Transfer
A key advantage of the RFP over previous works is the ability to fit novel and unseen grippers at test time via the unified coordinate map, e.g. from a human hand model. We evaluated grasp transfer from humans by collecting human grasp demonstrations and executing the transferred grasps with a Fetch robot gripper, shown in Fig. 7. For a human demonstration, we infer the MANO [33] hand parameters with the HaMeR [34] model. Using the sphere coordinate mapping defined in Sec. III-B, we have the correspondences for human model to the Fetch gripper used for grasp transfer. We evaluated the grasp transfer pipeline in real world on 14 YCB objects from SceneReplica [16] by making the Fetch robot grasp and lift the object using the transfer method. We achieve a grasping success rate of $0.785$ ($11/14$), with examples shown in Fig. 11. Most failures stem from an incorrect initialization from noisy human hand pose inference.
VII Conclusion and Future Work